### v2

# Token Embeddings Violate the Manifold Hypothesis

Abstract.
A full understanding of the behavior of a large language model (LLM) requires our understanding of its input token space. If this space differs from our assumptions, our understanding of and conclusions about the LLM will likely be flawed. We elucidate the structure of the token embeddings both empirically and theoretically. We present a novel statistical test assuming that the neighborhood around each token has a relatively flat and smooth structure as the null hypothesis. Failing to reject the null is uninformative, but rejecting it at a specific token $\psi$ implies an irregularity in the token subspace in a $\psi$-neighborhood, $B(\psi)$. The structure assumed in the null is a generalization of a manifold with boundary called a smooth fiber bundle (which can be split into two spatial regimes – small and large radius), so we denote our new hypothesis test as the “fiber bundle hypothesis.” Failure to reject the null hypothesis is uninformative, but rejecting it at $\psi$ indicates a statistically significant irregularity at $B(\psi)$. By running our test over several open-source LLMs, each with unique token embeddings, we find that the null is frequently rejected, and so the evidence suggests that the token subspace is not a fiber bundle and hence also not a manifold. As a consequence of our findings, when an LLM is presented with two semantically equivalent prompts, if one prompt contains a token implicated by our test, the response to that prompt will likely exhibit less stability than the other.
1. Introduction
A core area for Artificial Intelligence (AI) research is to understand the central components of modern learning systems, e.g. large language models (LLMs), assuming certain regularity conditions. Extensive research efforts have been mobilized to tackle the explainability of AI question ([1]), leveraging ideas from computer science ([2, 3, 4]), information theory ([5, 6]), and theoretical physics ([7, 8, 9]), to mathematical insights from high-dimensional probability theory ([10, 11]), and algebraic and differential topology ([12, 13, 14]) either to prove mathematical guarantees or to elucidate model behavior.
In many—if not most—AI research papers, a manifold hypothesis is tacitly assumed, that the data are concentrated near a low curvature manifold without boundary. This assumption drives the research just as an assumption of Gaussianity would drive data modeling. It is important to test this kind of assumption before data modeling. One common strategy to perform such a test is simply “plotting the data” to ascertain the empirical distribution. Plotting high-dimensional data, however, is non-trivial, but there are statistical tests to help us understand the underlying distribution (Mann-Whitney, Kolomogorov-Smirnoff, etc.). In this paper, we provide an analogous test for the regularity of token embeddings.
The first layer of most LLMs is usually comprised of the token embeddings, which are vector representations of tokens (fragments of human language, e.g. parts of words, or numbers). Thus, a good first step to explain the behavior of an LLM is to understand its token subspace where these embedding vectors live. For the remainder of this paper, the term “token”’ used in the context of the token subspace will refer to this embedding. Numerous papers (for instance, see [15, 16, 17, 18, 19, 20, 21]) have pointed to unexpected behaviors exhibited by LLMs that hinge on subtle changes in wording and text layout between apparently similar prompts, suggesting that certain semantically similar tokens may have dramatically different neighborhoods in the token subspace. These differences in neighborhoods correspond to places where the token subspace does not behave stably as it is irregular at such a token. Linguistically, these irregularities may correspond to polysemy or homonyms—tokens with multiple distinct meanings ([22]).
The presence of irregularities imply small changes in language can drive large variation in the token subspace. Moreover, if the token subspace of an LLM is, in fact, singular (not a manifold), irregularities can persist into the output of the LLM, perhaps unavoidably and potentially even because of its architecture. This is because the input and output of LLMs are a fixed finite set of tokens yet the internal architectures of the model assume a structure resembling a Riemannian manifold: a metrizable space that locally preserves inner products. Not accounting for singularities in the token subspace may confound our understanding of the LLM’s behavior.
1.1. Contributions
We present three hypothesis tests that determine for a given token $\psi$, whether its token subspace neighborhood $B(\psi)$ contains an irregularity, i.e. either has significantly high curvature or is singular. The tests are based on Theorem 1 that allows us to identify changes in subspace dimension that are inconsistent with the token subspace being:
  • 1.
    A manifold (without boundary) with a given embedded reach (low reach corresponds to high curvature and vice versa ([23])),
  • 2.
    A fibered manifold with boundary (which we call a fiber bundle for brevity) with the same embedded reach, estimated using small radii, and
  • 3.
    A fiber bundle with the same embedded reach, estimated using large radii.
Figure 1. This figure shows the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B. (The fiber bundle tests are not shown due to space considerations.) Since the latent space of Mistral7B has dimension $\ell=4096$, we have projected the coordinates of each token shown to $3$ using principal components analysis (PCA). The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region. Each point shown in this figure is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test. Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature. Putative geometric structures are labeled for certain tokens.
When these tests are rejected, this implies that $\psi$ is less stable under mathematical transformations than other tokens. In Section 3.2, we show the endemic nature of these instabilities by proving that context cannot resolve singularities persistently. And finally, we provide conditions (Theorem 2) for when singularities can propagate into the output of an LLM for generic transformers.
In our work, we provide empirical evidence that the global token embedding obtained from four different LLMs cannot be a manifold due to non-constant local dimension. Because our hypothesis tests require the distances between all pairs of tokens, we need the token embedding vectors. For this reason, we conduct our analysis to four open source models of moderate size – GPT2 ([24]), Llemma7B ([25]), Mistral7B ([26]), and Pythia6.9B ([27]). We emphasize that while this paper demonstrates the manifold and fiber bundle tests on four open source LLMs, our main contribution is primarily a methodology for hypothesis testing on embedded points. In particular, our method can be applied to any point cloud data set; the token embeddings we study here are merely one example.
Figure 2. (Top) Flowchart of our statistical methodology. (Bottom) The distribution of local dimensions estimated near tokens in GPT2, produced by our method as an intermediate step. This histogram shows a mixture with at least three distinct peaks.
1.2. Background and related work
Manifold learning takes its roots in nonlinear dimensionality reduction of high dimensional data into low dimensional representations of smooth manifolds. While these algorithms have played an important role in modern data science especially in domains such as cosmology ([28]), astrophysics ([29, 30]), dynamical systems ([31, 32]), and quantum chemistry ([33]), the main utility of these methods have been for data visualization. For instance, even though the Uniform Manifold Approximation and Projection (UMAP) provided an optimal algorithm for fitting sampled data to a manifold, the authors themselves concede that “if the global structure is of primary interest then UMAP may not be the best choice for dimensionality reduction” ([34]). These manifold learning algorithms will find some smooth structure to fit the data rather than “error out” when the data is not sampled from a manifold as it is not a hypothesis test. While UMAP will fit a manifold to data optimally, for instance, the manifold it finds may have extremely high curvature or dimension. Both of these are undesirable properties in a model, and both implicitly suggest that a manifold may be a poor choice for a representation. To address this concern, [35] presented a test for whether data lie on a manifold with a specified embedded reach. Unfortunately, [35] does not include the values of key constants, which makes the practical usage of this test impossible. Our paper remedies both issues: we construct a practical hypothesis test for manifolds and fiber bundles with a specified embedded reach and implement it as software.
In our test case, the basic assumption is that transformers act on the entire extrinsic latent space containing the token embedding. This is clearly a manifold as it is Euclidean. But the truth is that a transformer in the context of an LLM really only acts on the embedded token subspace. The token input embedding matrix defines this subspace, by specifying where the tokens are located. Such a space is quite unconstrained. There is no a priori reason to suspect it is a manifold. It has already been shown that local neighborhoods of each token have salient topological structure ([13]), and one of the most basic parameters is the dimension near any given token in this space. Higher dimension at a token $\psi$ means that $\psi$ has greater density while lower dimensional tokens have less density ([22]). Intuitively, a densely populated $B(\psi)$ will generically occupy a higher dimension than otherwise.
Assuming that word embeddings yield manifolds, some researchers have used global dimension estimators on token input embeddings and word embeddings ([36, 37]). By using a local (not global) dimension estimator, [38] presented the first (to our knowledge) direct test of whether the token subspace is a manifold for the token input embeddings for several LLMs, though no rigorous hypothesis test was presented. Nevertheless, a strongly negative result was obtained: the subspace of tokens is apparently never a manifold, so global dimension estimators are not reliable.
2. Methods
Our method, summarized in Figures 2 and 3, proposes three related hypotheses: (1) the manifold hypothesis, (2) the fiber bundle hypothesis for small radius, and (3) the fiber bundle hypothesis for larger radius.
Figure 3. (a) Schematic of our three hypothesis tests (b) Application of the tests to the tokens #, ¢, and \$ with GPT2’s token input embedding.
The key insight is that Theorem 1 asserts that for a fiber bundle, the number of tokens within a ball of a given radius $r$ of a fixed token is a piecewise log-log-linear function of radius, in which the slopes correspond to dimension and decrease as $r$ increases. If there is a statistically significant change in the slope estimates (here, the significance level $\alpha=10^{-3}$), then we reject the manifold hypothesis. If there is a statistically significant increase in slope, then we additionally reject the fiber bundle hypothesis in the corresponding region (small or large radius).
Figure 3(b) demonstrates our method on three tokens used by GPT2. The tokens \$ and # play an important role in many programming languages, whereas ¢ does not. (Note that these tokens were chosen for illustrative purposes. The $p$-value for rejecting the hypotheses is larger than $\alpha=10^{-3}$ used in Section 3.3, so these particular tokens do not appear in any of the tables in the following sections.) The token # does not show any rejections of our hypotheses. While this does not allow one to conclude that the vicinity of # is a manifold or fiber bundle, we can use Theorem 1 to estimate the dimension from the slope of the curve (approximately $53$, once logarithms of the values on both axes are taken). The token ¢ exhibits a rejection of the manifold hypothesis, but not the fiber bundle hypothesis. We can conclude that there is a irregularity at a distance of roughly $3.9$ units of ¢. The curve for \$ exhibits two slope changes at radii $3.8$ and $4.0$. The smaller radius slope change represents a violation of the manifold hypothesis, while the large radius slope change represents a violation of the fiber bundle hypothesis.
3. Results
3.1. Volume versus radius for fiber bundles
The learned token embedding vectors in an LLM each have the same number of elements, which we refer to as $\ell$. Thus, the explicit representation of the token subspace $T$ used by an LLM arises by embedding $T$ within the latent space $\mathbb{R}^{\ell}$ by way of a smooth function $e:T\to\mathbb{R}^{\ell}$. Since embeddings are bijective, each $\psi\in T$ corresponds to a unique element $e(\psi)\in\mathbb{R}^{\ell}$. With a slight abuse of notation, we can write $\psi\in e(T)$.
The fiber bundle model imposes a specific assumption about the structure on $T$, by relating it to another, lower dimensional, manifold $S$, called the base space, via a smooth map $p:T\to S$.
Definition 1.
A fibered manifold is a surjective function $p:T\to S$ such that the Jacobian matrix $J_{\psi}p$ at every point $\psi\in T$ has rank equal to the dimension of $S$. For brevity, we refer to this kind of space simply as a fiber bundle.
If a fiber bundle is the correct representation of the space of tokens $T$, then the probability distribution of tokens given by the volume form $v$ on $T$ will impact the distribution of tokens within $\mathbb{R}^{\ell}$.
Theorem 1.
Suppose that $T$ is a compact, finite-dimensional Riemannian manifold with boundary [^1] Every point in a manifold with boundary has a neighborhood that is locally homeomorphic to a half-space. As a consequence, manifolds are a special case of manifolds with boundary., with a volume form $v$ satisfying $v(T)<\infty$, and let $p:T\to S$ be a fibered manifold. If $e:T\to\mathbb{R}^{\ell}$ is a smooth embedding with reach $\tau$, then there is a function $\rho:e(T)\to[0,\tau]$ such that if $\psi\in e(T)$, the induced volume $(e_{*}v)$ in $\mathbb{R}^{\ell}$ satisfies
(1) $(e_{*}v)\left(B_{r}(\psi\right)=\begin{cases}O\left(r^{\dim T}\right)&\text{if}0\leq r\leq\rho(\psi),\\ (e_{*}v)\left(B_{\rho(\psi)}(\psi)\right)+O\left((r-\rho(\psi))^{\dim S}\right)&\text{if }\rho(\psi)\leq r,\\ \end{cases}$
where $B_{r}(\psi)$ is the ball of radius $r$ centered at $\psi$, and the asymptotic limits are valid for small $r$.
Proof.
See Section A.4
Theorem 1 characterizes the resulting probability distribution using parameters (the exponents in Equation (1)) that can be estimated from the token input embedding. These parameters are bounded by the dimensions of $S$ and $T$.
In short, if we take the logarithm of both the volume and radius in Equation (1), we obtain a piecewise linear curve, in which the slopes are given by $\dim S<\dim T$, and the slopes must decrease through a discontinuity as the radius increases, as shown in Figure 3(a).
3.2. Context does not persistently resolve singularities
Given a context window of $w-1$ tokens, an LLM uses its learned parameters to predict the next ($w$-th) token. This gets added to the context window and the process continues. (See Section A.1 for details.) In this section, we provide a mathematical argument showing that the use of context is insufficient to resolve singularities present in the token subspace for a generic transformer in a persistent manner. Again, denote the token subspace by $T$, noting that Section 3.3 shows that $T$ is unlikely to be a manifold for typical LLMs. Let us begin with a lemma:
Lemma 1.
If $X$ is a manifold, $Y$ is an arbitrary topological space, and $X\times Y$ is a manifold, then $Y$ must also be a manifold.
Proof.
Observe that $Y$ is homeomorphic to $(X\times Y)/X$, which is the quotient of two manifolds via the projection map, which is a surjective submersion. Hence $Y$ is a manifold. ∎
Consequently, the contrapositive of Lemma 1 states that if $T$ is not a manifold but $T^{w-1}$ is a manifold, it follows that the space of sequences of tokens [^2] Caution: $T^{w}$ is the sequence of tokens embedded individually, which is different from the embedding of the entire sequence. Theorem 2 asserts conditions under which $T^{w}$ has a direct impact on LLM output. $T^{w}$ cannot be a manifold. At best, the latent space for every other context window size can be a manifold. For every proposed context window length, a longer window will exhibit singularities. In short, lengthening the context window does not persistently resolve singularities in the token embedding.
The remaining question is whether these singularities persist into the output of the LLM. Can the user elicit the instabilities caused by singularities? We provide sufficient conditions for unstable outputs:
Theorem 2.
Let $Z$ be the $d$-dimensional bounding manifold [^3] Since $T\subseteq\mathbb{R}^{\ell}$, it is always contained within such a bounding manifold $Z\subseteq\mathbb{R}^{\ell}$. for the token subspace, such that $T\subseteq Z$. Consider an LLM with a context window of size $w$, in which the latent space of tokens is $\mathbb{R}^{\ell}$, and we collect $m$ tokens as output from this LLM.
Suppose the following:
  1. (1)
    (enough tokens are collected from the response) $m>\frac{2wd}{\ell}$, but
  2. (2)
    (the context window is longer than the number of tokens we collected) $w\geq m$.
Under these conditions, a generic set of transformers yield a topological embedding of $T^{w}$ into the output of the LLM.
Proof.
See Section A.4
As a consequence, there always exists such an $m$ satisfying the hypothesis if $2d<\ell$, i.e. the bounding manifold has dimension less than half the latent space dimension. If our hypothesis tests show that the token subspace contains singularities and the hypotheses of Theorem 2 are satisfied, singularities will persist into the output of the LLM for generic transformers even for large context window sizes.
3.3. Experimental results
We applied our test to the token subspaces of the four open source LLMs introduced in Sec. 4.1: GPT2, Llemma7B, Mistral7B, and Pythia6.9B. In each LLM we tested, we found evidence that the token subspace is not globally a manifold because the local dimension is highly variable. Moreover, we can infer that they are either not fiber bundles or have high curvature locally; in either case—violating the hypotheses or having high curvature—model behavior near certain tokens will be highly unstable.
We observe highly statistically significant differences in the irregular tokens between LLMs—even for those with identical sets of tokens overall—which indicates that their respective training methodologies have a strong impact on the token subspace. Given these differences, prompts containing irregular tokens will likely produce dissimilar outputs across the four LLMs corroborating the results shown by [17]. Additionally, prompts containing irregular tokens within an LLM will likely be unstable and produce highly variable outputs with replicated queries. This is clear if $B(\psi)$ contains a singularity but will also be true for regions of high curvature resulting in instability for both interpolation and extrapolation tasks ([39]).
Several non-manifold features are visibly present in Figure 1. The tokens “ember” and “P” appear to be both putative examples of cusp points, places where the space exhibits a sharp exposed point. Manifolds cannot have cusp points, so attempting to approximate a cusp with a manifold—for instance, with UMAP—necessitates high curvature. The tokens “:” and “ly” appear to be pinch points—like what are present in a chain of sausages—which again cannot be approximated by a manifold unless it has high curvature.
Several additional views of Figure 1 are shown in Figure 6 in the Supplement. The coordinates, token text, $p$-values, and a video of the space with commentary are available in the ZIP file Supplement.
Table 1. Dimensional data for and number of tokens rejecting the manifold and fiber bundle hypotheses at two slope changes. $n$ is the number of tokens (vocabulary size) in each model. The $p$-values displayed report the minimum $p$-value for each test, accounting for multiple testing. Quartiles (Q2 is the median) are shown for the distribution of the dimensions estimated from the slopes of the log-volume versus log-radius plots aggregated over all tokens. The “rejects” columns list the number of tokens rejecting each hypothesis. Table 3 (resp. Table 49) in the Supplementary material shows each of the fiber bundle (manifold hypothesis) violations for each model that is listed here.
Model Manifold Fiber bundle
rejects Smaller Radius Larger radius
dim. rejects dim. rejects
GPT2 $68$ Q1: $20$ $12$ Q1: $8$ $7$
$n=50257$ $p\approx 3\times 10^{-8}$ Q2: $389$ $p\approx 9\times 10^{-6}$ Q2: $14$ $p\approx 3\times 10^{-8}$
Q3: $531$ Q3: $32$
Llemma7B $33$ Q1: $4096$ 00 Q1: $8$ $1$
$n=32016$ $p\approx 5\times 10^{-9}$ Q2: $4096$ N/A Q2: $11$ $p\approx 3\times 10^{-4}$
Q3: $4096$ Q3: $14$
Mistral7B $40$ Q1: $9$ $1$ Q1: $5$ $2$
$n=32016$ $p\approx 3\times 10^{-7}$ Q2: $48$ $p\approx 8\times 10^{-4}$ Q2: $6$ $p\approx 8\times 10^{-5}$
Q3: $220$ Q3: $9$
Pythia6.9B $54$ Q1: $2$ 00 Q1: $2$ 00
$n=50254$ $p\approx 2\times 10^{-7}$ Q2: $108$ N/A Q2: $5$ N/A
Q3: $235$ Q3: $145$
Table 1 shows the results of our hypothesis tests for the four models we analyzed. It is clear that the models have quite different token input embeddings, and all of them exhibit highly significant rejections of the manifold hypothesis. GPT2, Llemma7B and Mistral7B also reject the fiber bundle hypothesis. The rejections of the fiber bundle hypothesis are more frequent in the larger radius region than the smaller radius region, which is consistent with the polysemy interpretation of [22].
While most of the tokens are not shared between the LLMs, Llemma7B and Mistral7B do have identical token sets. The fact that Table 1 shows significant differences between these two models indicates that the structure of the irregularities for these these two models is quite different. This implies that their response to the same prompt is expected to be markedly different, even without considering the differences in their respective transformer stages.
Since there are many rejections of the manifold hypothesis, we summarize some general trends.
  • The GPT2 tokens at irregularities are tokens that can only appear at the beginning of words.
  • The Pythia6.9B tokens at irregularities are nearly all word fragments or short sequences of text that are quite meaningless on their own.
  • The Llemma7B and Mistral7B tokens at irregularities are a combination of the previous two: either they can only appear at the beginning of words or they are word fragments.
Table 2. Testing the hypotheses of Theorem 2: Will singularities generically persist into the output? See statement of Theorem 2 for variable definitions. “Small” and “Large” refer to the small and large radius estimates of dimensions from Table 1.
Model Latent dim Bounding dim. Context Min. output tokens Satisfied?
$\ell$ $d$ $w$ $m$ such that $w\geq m$?
Small Large Small Large Small Large
GPT2 $768$ $389$ $14$ $1024$ $1038$ $38$ No Yes
Llemma7B $4096$ $4096$ $11$ $4096$ $8193$ $23$ No Yes
Mistral7B $4096$ $48$ $6$ $4096$ $97$ $13$ Yes Yes
Pythia6.9B $4096$ $108$ $5$ $4096$ $217$ $11$ Yes Yes
Estimates of the bounding manifold dimension used in Theorem 2 for the four LLMs we studied are in the “dimension” columns of Table 1, with the most stringent being the dimension for smaller radius. Table 2 shows the results of testing the hypotheses of Theorem 2 using the medians (Q2). Mistral7B, and Pythia6.9B satisfy the hypotheses of Theorem 2 using the small radius dimension estimates, but GPT2 and Llemma7B do not. All four LLMs satisfy the hypotheses of Theorem 2 if we instead use the larger radius dimension estimates. This means that it is likely that the transformers of Mistral7B, and Pythia6.9B are unable to resolve the singularities that we show exist in the token subspace, so singularities will persist in the output of the model. If the true bounding manifold for GPT2 or Llemma7B is smaller than the small radius estimate suggests, these too will likely produce singularities in their output.
4. Discussion
None of the four LLMs we studied are manifold globally given non-constant local dimension. Locally, the token subspace cannot be manifolds with large reach, and three of the four also cannot be fiber bundles with large reach. Irregularities—tokens that cause rejections of the manifold hypothesis—occur in different ways across all four LLMs. Additionally, irregularities corresponding to violations of the fiber bundle hypothesis are tokens whose neighborhoods exhibit a dependency between the large- and small-scale variability.
The grammatical role of tokens is likely a root cause for some of sensitivity of LLMs to their prompts that has been observed in the literature ([20, 21]), and may explain why “explaining LLM behavior” remains challenging and open. Most methods for explaining LLM behavior in terms of dynamical systems, for instance, derive their inferential power from assuming that the token subspace is a manifold ([40, 41, 42]). Our results indicate that these theoretical methods may not apply to current LLMs.
One might argue that context helps to disambiguate and “fix” these problems, but this intuitive argument seems not always possible as context alone is insufficient to resolve singularities. As we showed in Section 3.2, singularities in the token subspace persistently lead to singularities in the space of token sequences, and under appropriate conditions these propagate into the output of an LLM.
Irregularities—which may in fact be singularities—are noticeable in a PCA projection of the token subspace (Figure 1 and Figure 6). Intuitively, this means that the use of inner products to perform token de-embeddings by orthogonal projections may not be accomplishing the intended objectives. Even small perturbations within the latent space may result in instabilities in these projections, providing potential explanations for the instabilities in LLM outputs.
In addition to token instability, the fact that token subspaces are not manifolds and may have small reach means that the geodesic distance between tokens can also be very unstable. While the distance along geodesics can be defined, it may not correlate with any sense of semantic distance between tokens. Furthermore, as [38] found, in most of the models, there are tokens with dimension 00 neighborhoods. These tokens are theoretically isolated, implying that the token subspace is disconnected and hence intrinsically infinitely far from other tokens.
Singularities may arise either as artifacts of the training process or from features of the languages being represented. Consistent with the idea that polysemy may yield singularities ([22]), several of the tokens where the fiber bundle hypothesis is rejected are clear homonyms. For instance, both “affect” and “monitor” can be used as either nouns or verbs, and their meanings are different in these two roles. Tokens also form fragments of text, and so, a token may correspond to homonyms after the addition of a prefix or suffix. A token like “aunder” can be prefixed to yield the word “launder”, a word with multiple meanings of opposite sense. Specifically, one can “launder” clothing (positive connotation) or “launder” money (negative connotation). Several other tokens where the fiber bundle hypothesis is violated form words with substantially different meanings or grammatical roles upon adding a prefix or suffix. For instance, “wins” can function as a noun, as a verb, or form the adjective “winsome”.
The differences in how the manifold and fiber bundle hypotheses are rejected across different LLMs suggest that the training methodology for each model leaves an indelible fingerprint (or gDNA coined by [17]). Making general assertions about LLMs without consideration of their training and evolution is likely fraught. Even between Llemma7B and Mistral7B, which have identical token sets, prompts likely cannot be “ported” from one LLM to another without significant modifications if they contain tokens near irregularities. This may have serious ramifications for interacting LLMs such as those powering multi-agentic frameworks being deployed today ([43, 44, 45]).
A few clear patterns among tokens near irregularities are nevertheless noticeable. Tokens that begin a word or are a word fragment are often located at an irregularity. Additionally, in Llemma7B (but not Mistral7B) and Pythia6.9B the tokens with unusually low fiber dimension often contain non-printing or whitespace characters. This suggests that these models are quite sensitive to text layout, perhaps to the exclusion of more semantically salient features in the text. Given our findings, future experiments can be run to explore the impact of irregular tokens on the variability of responses produced by different LLMs.
As we attempt to model natural phenomena particularly in our physical world, we know singularities exist (e.g. phase transitions), and so assuming regularity limits our understanding of the natural world. Our work extends this notion to the linguistics of natural language where its richness and complexity do not appear to manifest in a smooth representation.
4.1. Limitations
While topological theory motivates our statistical test, only the slope changes that occur for radii less than the embedded reach correspond to rejections of the manifold or fiber bundle hypotheses. If the reach of a manifold is small, then it likely has a high local curvature, and measuring distances on that surface will be numerically unstable. As a result, the reach itself is difficult to estimate accurately from sampled data ([23, 46, 47]). Our implementation tests the radius below an estimated reach, determined by an inspection of the radius-vs-volume plots of a random sample of tokens. If the true reach is smaller than our estimates, the curvature of the manifolds that fit the data well, found by any method, will be high.
And lastly, we emphasize that these tests allow us to reject the structure assumed in our null hypothesis (smooth and regular), but do not permit us to infer an alternative. While rejecting regularity does imply singularity, it does not afford us any evidence as to the type of singularity. That must be determined post hoc if possible.
Acknowledgments
The authors would like to thank Mohammed Abouzaid, Anand Sarwate, Andrew Lauziere, and Andrew Engel for helpful suggestions on a draft of this manuscript.
This material is based upon work partially supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001124C0319. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA). Distribution Statement “A” (Approved for Public Release, Distribution Unlimited) applies to the portion of the work funded by DARPA.
Appendix A Supplementary Material
A.1. Background
At an abstract but precise level, an LLM consists of several interacting processes, as outlined in Figure 4. An LLM implements a transformation of a sequence of tokens (the query) into a new sequence of tokens (the response). Formally, if each input token is an element of a metric space $T$, then the LLM is a transformation $T^{w}\to T^{m}$, where $w$ is the number of tokens in the query and $m$ is the number of tokens in the response. This transformation is typically not a function because it is stochastic—it involves random draws.
Figure 4. Data flow in a typical LLM. A sequence of tokens forming the query is converted via the token input embedding $e^{w}$ into the initial context window, as a point in the latent space $X^{w}$. Each of these windows in the latent space are converted, token-by-token, into probability distributions via $f$ into the single token latent space $X$. From these, each token presented in the output (in the set $Y$) is obtained via a random draw. These output tokens are then used for subsequent windows.
To operate upon tokens using numerical models, such as could be implemented using neural networks, we must transform the finite set of tokens $T$ into numerical data. This is typically done by way of a pair of latent spaces $X=\mathbb{R}^{\ell}$ and $Y=\mathbb{R}^{q}$. The dimension $q$ of $Y$ is chosen to be equal to the number of elements in $T$, so that elements of $Y$ have the interpretation of being (unnormalized) probability distributions over $T$.
The transformation $T^{w}\to T^{m}$ is constructed in several stages.
Input tokenization:
: Each token is embedded individually via the token input embedding function $e:T\to X$. As a whole, $X^{w}$ is called a latent window.
Transformer blocks:
: The probability distribution for the next token is constructed by a continuous function $f:X^{w}\to Y$. This is usually implemented by one or more transformer blocks.
Output tokenization:
: Given the output of one of the transformer blocks $f$, one can obtain an output token in $T$ by a random draw. Specifically, if $(\psi_{1},\psi_{2},\dotsc,\psi_{w})$ is the current window in $X^{w}$, then the next token $t$ is drawn from the distribution given by $f(\psi_{1},\psi_{2},\dotsc,\psi_{w})$.
Next window prediction:
: Given that token $t$ was drawn from the distribution, the next latent window itself is constructed by a transformation $F:X^{w}\to X^{w}$, which advances the window as follows:
$F(\psi_{1},\psi_{2},\dotsc,\psi_{w}):=(\psi_{2},\dotsc,\psi_{w},t).$
Note well: the function $e^{w}:T^{w}\to X^{w}$ which transforms a sequence of individual tokens into coordinates in the latent space is merely the first stage in the process. As such, $e^{w}$ is distinct from the embedding of sequences provided by many LLM software interfaces. That sequence embedding is properly an intermediate stage within the function $f$.
The focus of this paper is specifically upon the structure of the token input embedding $e:T\to X=\mathbb{R}^{\ell}$. Since the token set $T$ is finite, $e$ can be stored as a matrix. In this matrix, each column corresponds to an element of $T$, and thereby ascribes a vector of numerical coordinates to each token. By replacing the last layer of the deep neural network $f:X^{w}\to Y$, a vector of probabilities for the next token is obtained from the activations of the last layer. One can therefore interpret the probabilities as specifying a token output embedding. Both the tokenization and the transformer stages are learned during training, and many strategies for this learning process are discussed extensively in the literature. These two stages interact when they produce the LLM output, so it is important to understand the lineage of a given tokenization as being from a particular LLM.
A.2. Computational resource requirements
Applying our entire method (starting with the token embedding matrix and ending with the $p$-values for each of the three tests at every token) to each model took approximately 12 hours of wall clock time (per model) on an Intel Core i7-3820 with 32 GB of CPU RAM and no GPU running at 3.60GHz. Almost all of the runtime was spent computing the token distance matrix. The subsequent stages, including the three hypothesis tests, shown in Figure 2 run in only a few minutes on the same machine. The distance matrix calculation was performed using scipy.spatial.distance_matrix(), which produces a dense matrix, by computing the distance between all pairs of tokens in the latent space. Because of the relatively low runtime requirements, we did not feel the need to use sparsity-enhancing techniques to accelerate the distance matrix calculation.
A.3. Interpretation of fiber bundles in terms of spatial variability
It is usual to describe measurements as exhibiting the combined effect of multiple spatial scales. One way this might arise is if we could express each measurement as being an ordered pair (signal, noise), with the short spatial scales correlating to noise and the longer spatial scales correlating to the signal. With this interpretation, if the space of all possible signals is $S$ and the space of all noise values is $V$, we could represent the space of all possible measurements as the cartesian product $E=S\times V$. This cartesian product representation can arise for noiseless data as well, as it merely captures multiple spatial scales. In what follows, we will call $S$ the base space and $V$ the fiber space.
Figure 5. Our method applied to a fiber bundle in $\mathbb{R}^{2}$. The vertical direction is the base space (long spatial scale), while the horizontal direction represents the fibers space (short spatial scale). Gray points on the right frame show estimates from a random sampling of points in the strip; the solid line shows the theoretical area versus radius curve.
Figure 5 shows an example of this situation. It consists of a $1$-dimensional base space (the long spatial scale) and $1$-dimensional fibers (the short spatial scale), which in this case forms a narrow strip in the plane. Volumes (areas, in this case) of balls of small radius scale quadratically (slope $2$ in a log-log plot), but scale asymptotically linearly (slope $1$ in a log-log plot) for large radii. The transition between these two behaviors is detectable by way of a corner in the plot. This situation is easily and robustly estimated from the data; the gray points in Figure 5 (right) are derived from a random sampling of points drawn from the strip.
To test for multiple spatial scales, we propose a signal model that is mathematically represented by a fiber bundle. In a fiber bundle, the long spatial scales are modeled by a space $S$, but all possible measurements are modeled by a function $p:E\to S$. The idea is that fibers $p^{-1}(b)$ are still cartesian products: pairs of both spatial scales, and these are all identical up to diffeomorphism. Our method relies upon a particular geometric property of fiber bundles: we can identify if the fibers are not all identical according to when the conclusion of Theorem 1 is violated.
One possible way that a fiber bundle might arise is if the underlying noise model is homooscedastic, which is to say that the dimensionality of the noise near a token does not depend on that token. In this case, the fibers (short spatial scales) mostly correspond to noise, and the base (long spatial scales) correspond mostly to signal. Nevertheless, a fiber bundle may arise with noiseless data as well; the presence of multiple spatial scales is what matters.
On the other hand, if the underlying noise model is heteroscedastic—it depends on the token in question—then the fiber bundle model should be rejected. Figure 3(a) shows a situation that is not a fiber bundle, since there is a change in the dimension of the fiber. In the upper portion of the figure, the fiber dimension is 00 while in the lower portion the fiber dimension is $1$. This is detectable by looking at the volume versus radius plots for two samples. While both samples show corners in their volume versus radius plots, Theorem 1 establishes that the slopes always decrease with increasing radius for a fiber bundle. This is violated for the sample marked $\psi_{1}$, so we conclude that the space is not a fiber bundle. On the other hand, because the sample marked $\psi_{2}$ does not exhibit this violation, it is important to note that if a sample yields data consistent with Theorem 1, we cannot conclude that the space is a fiber bundle.
A.4. Mathematical proof of Theorems 1 and 2
This section contains mathematical justification for the fiber bundle hypothesis proposed earlier in the paper and the proof of Theorems 1 and 2. The central idea is the use of a special kind of fiber bundle, namely a fibered manifold with boundary.
By the submersion theorem ([48]), if the Jacobian matrix of a fibered manifold $p:T\to S$ at every point has rank equal to the dimension of $S$, then the preimages $p^{-1}(\psi)\subseteq T$ of each point $\psi\in T$ are all diffeomorphic to each other. These preimages form the fibers discussed in the earlier sections of the paper.
As a consequence, each point $y$ in the base space $S$ has an open neighborhood $U$ where the preimage $p^{-1}(U)$ is diffeomorphic to the product $U\times p^{-1}(y)$, which is precisely the base-fiber split discussed in Section A.3. Specifically, the base dimension is simply the dimension of $S$, whereas the fiber dimension is the $(\dim T-\dim S)$.
The notion of a fibered manifold $p:T\to S$ forms the intrinsic model of the data, which is only implicit in an LLM. The tokens present in a given LLM can be thought of as a sample from a probability distribution $v$ on $T$, which can be taken to be the Riemannian volume form on $T$ normalized so that $v(T)=1$.
Definition 2.
If $e:T\to\mathbb{R}^{\ell}$ is a continuous map and $v$ is a volume form on $T$, then the pushforward is defined by
$(e_{*}v)(V):=v\left(e^{-1}(V)\right)$
for each measurable set $V$.
It is a standard fact that if $e$ is a fibered manifold or an embedding, then $e_{*}v$ is also a volume form.
Proof.
(of Theorem 1) If $e$ is assumed to be a smooth embedding [^4] Smoothness is not a strong constraint here. Any continuous function (such as ReLU) can be approximated well enough by a smooth map., the image of $e$ is a manifold of dimension $\dim T$. The pushforward of a volume form is a contravariant functor, so this means that $e_{*}v$ is the volume form for a Riemannian metric on $e(T)$. Using this Riemannian metric on $e(T)$, then [49, Thm 3.1] implies that for every $\psi\in e(T)$, if $r\ll\tau$, then
(2) $(e_{*}v)\left(B_{r}(\psi)\right)=O\left(r^{\dim T}\right).$
Since $T$ is compact, $S$ is also compact via the surjectivity of $p$. This implies that there is a maximum radius $r_{1}$ for which a ball of this radius centered on a point on $\psi\in e(T)$ is entirely contained within $e(T)$. Also by compactness of $S$, there is a minimum radius $r_{2}$ such that a ball of radius $r_{2}$ centered on a point $\psi\in e(T)$ contains a point outside of $e(T)$.
Since $e$ is assumed to be an embedding, by the tubular neighborhood theorem ([48]), it must be that $r_{2}<\tau$. Define
$\rho(\psi):=\textrm{argmax }_{r}\left\{B_{r}(\psi)\subseteq e(T)\right\},$
from which it follows that $0<r_{1}\leq\rho(\psi)\leq r_{2}<\tau$. As a result, Equation (2) holds for all $r\leq\rho(\psi)$, which is also the first case listed in Equation (1).
If $r$ is chosen such that $\rho(\psi)<r<\tau$, the volume of the ball centered on $\psi$ of radius $r$ will be less than what is given by Equation (2), namely
$(e_{*}v)\left(B_{r}(\psi)\right)<O(r^{\dim T}).$
Since $v$ is a volume form, its pushforward $(p_{*}v)$ onto $S$ is also a volume form. Moreover, via the surjectivity of $p$,
$(e_{*}v)(B_{r}(\psi))$ $=v(e^{-1}(B_{r}(\psi)))$
$\leq v(p^{-1}(p(e^{-1}(B_{r}(\psi)))))$
$\leq(p_{*}v)(p(e^{-1}(B_{r}(\psi))))$
$\leq O(r^{\dim S}).$
From this, the second case of Equation (1) follows by recentering the asymptotic series on $\rho(\psi)$. ∎
Notice that the second case in Equation (1) may be precluded since while it holds for small $r$, it may be that $\rho(\psi)$ may not be sufficiently small. As a consequence, the second case only occurs when both $r$ and $\rho(\psi)$ are sufficiently small. In the results shown in Section 3.3, both cases appear to hold frequently.
We now address Theorem 2.
Proof.
(of Theorem 2) We need to manipulate the inequalities in the hypothesis to show that they satisfy [50, Thm. 1]. Once the hypothesis is satisfied, [50, Thm. 1] establishes that a generic set of transformers will embed $T^{w}$ into the output of the LLM. Notice that [50, Thm. 1] considers the response of the LLM to a single token, though this can be easily extended by enlarging the bounding manifold dimension. With this in mind, the hypothesis to be satisfied is that
$2wd<m\ell\leq w\ell.$
The right inequality is clearly satisfied if context window is longer than number of output tokens collected, namely $w\geq m$. The left inequality is satisfied if
$2d<\frac{m}{w}\ell,$
which can be easily rearranged to the form in the hypothesis. ∎
A.5. Additional views for Figure 1
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6. Additional rotational views of the PCA of the neighborhood of ember in Mistral7B shown in Figure 1. Each point shown is a token. Colors are $p$-values for the manifold hypothesis, with darker colors corresponding to lower likelihood that that point’s neighborhood is a manifold with low curvature.
See Fig. 6.
A.6. Table of fiber bundle rejections in all four LLMs
Table 3. Violations of the fiber bundle hypothesis
Model Token Slope change $p$-value Comment
GPT2 laughable Smaller radius $9\times 10^{-6}$ Must start a word
GPT2 nuance Smaller radius $2\times 10^{-4}$ Must start a word
GPT2 dt Smaller radius $2\times 10^{-4}$
GPT2 Mesh Smaller radius $2\times 10^{-4}$
GPT2 affect Smaller radius $3\times 10^{-4}$ Must start a word
GPT2 Thankfully Smaller radius $3\times 10^{-4}$
GPT2 swat Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 Malaysian Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 Palestinian Smaller radius $7\times 10^{-4}$ Must start a word
GPT2 wins Smaller radius $8\times 10^{-4}$ Must start a word
GPT2 hedon Smaller radius $9\times 10^{-4}$
GPT2 donor Smaller radius $9\times 10^{-4}$ Must start a word
GPT2 Xan Larger radius $3\times 10^{-8}$ Must start a word
GPT2 aunder Larger radius $2\times 10^{-4}$
GPT2 Dri Larger radius $2\times 10^{-4}$
GPT2 ney Larger radius $3\times 10^{-4}$
GPT2 rodu Larger radius $3\times 10^{-4}$
GPT2 Insert Larger radius $4\times 10^{-4}$
GPT2 Ying Larger radius $4\times 10^{-4}$ Must start a word
Llemma7B pax Larger radius $3\times 10^{-4}$
Mistral7B änge Smaller radius $8\times 10^{-4}$
Mistral7B H0 Larger radius $5\times 10^{-4}$
Mistral7B monitor Larger radius $8\times 10^{-5}$ Must start a word
See Table 3.
A.7. Tables of manifold rejections in all four LLMs
Table 4. Rejections of the manifold hypothesis for GPT2 (1 of 2)
Number Token $p$-value Comment
1 Xan 0.0000000280 Must start a word
2 chemist 0.0000000602
3 fry 0.00000564 Must start a word
4 laughable 0.00000807 Must start a word
5 Retro 0.0000108 Must start a word
6 laughable 0.0000156 Must start a word
7 Thankfully 0.0000339
8 intoler 0.0000514 Must start a word
9 assumes 0.0000533 Must start a word
10 handwritten 0.0000783 Must start a word
11 MY 0.0000811 Must start a word
12 nuance 0.000105 Must start a word
13 playground 0.000110 Must start a word
14 dt 0.000124
15 aunder 0.000132
16 Mesh 0.000135
17 Disney 0.000139
18 McCl 0.000155 Must start a word
19 Dri 0.000196
20 Destiny 0.000206 Must start a word
21 ney 0.000206
22 juvenile 0.000209 Must start a word
23 advent 0.000211 Must start a word
24 perpet 0.000213 Must start a word
25 otally 0.000223
26 affect 0.000226 Must start a word
27 efficient 0.000238
28 rodu 0.000242
29 Kl 0.000260 Must start a word
30 Thankfully 0.000278
31 duct 0.000284
32 frail 0.000292 Must start a word
33 alle 0.000304 Must start a word
34 Cha 0.000322 Must start a word
35 OO 0.000333
Table 5. Rejections of the manifold hypothesis for GPT2 (2 of 2)
Number Token $p$-value Comment
36 Insert 0.000336
37 Ying 0.000357 Must start a word
38 Georgia 0.000373
39 flo 0.000384
40 uddled 0.000413
41 subsequ 0.000421 Must start a word
42 lig 0.000439 Must start a word
43 Environmental 0.000456 Must start a word
44 Vladimir 0.000466 Must start a word
45 Defence 0.000510 Must start a word
46 swat 0.000558 Must start a word
47 registering 0.000564 Must start a word
48 Malaysian 0.000569 Must start a word
49 trough 0.000575 Must start a word
50 ears 0.000615 Must start a word
51 Meh 0.000637 Must start a word
52 Palestinian 0.000648 Must start a word
53 reckless 0.000654 Must start a word
54 Wax 0.000658 Must start a word
55 Dist 0.000673 Must start a word
56 Quad 0.000725 Must start a word
57 relations 0.000749
58 wins 0.000781 Must start a word
59 KER 0.000785
60 gist 0.000786 Must start a word
61 hedon 0.000834
62 MMA 0.000839 Must start a word
63 donor 0.000865 Must start a word
64 aleb 0.000873
65 Asked 0.000873
66 Isaiah 0.000940 Must start a word
67 Survey 0.000947 Must start a word
68 youngster 0.000987 Must start a word
Table 6. Rejections of the manifold hypothesis for Llemma7B
Number Token $p$-value Comment
1 clojure 0.00000000434
2 agu 0.000000846
3 ¡Cyrillic¿ 0.000000938 Must start a word
4 custom 0.00000163
5 kW 0.00000496 Must start a word
[…]
7 porque 0.0000128 Must start a word
8 Leopold 0.0000135 Must start a word
9 Rem 0.0000232 Must start a word
10 Ern 0.0000265 Must start a word
11 though 0.0000540 Must start a word
12 č 0.0000619 Must start a word
13 endl 0.0000683 Must start a word
14 corrected 0.0000738 Must start a word
15 tw 0.0000788 Must start a word
16 ères 0.000115
17 grep 0.000151
18 kind 0.000202 Must start a word
19 pax 0.000243
20 Tri 0.000250 Must start a word
21 spielte 0.000258 Must start a word
22 Heaven 0.000262 Must start a word
23 d 0.000289
24 ¡Cyrillic¿ 0.000320
25 Card 0.000335 Must start a word
26 Wall 0.000351 Must start a word
27 periodic 0.000359 Must start a word
28 ucht 0.000463
29 jeu 0.000553 Must start a word
30 slightly 0.000807 Must start a word
31 passion 0.000823 Must start a word
32 Charles 0.000833 Must start a word
33 victory 0.000906 Must start a word
Table 7. Rejections of the manifold hypothesis for Mistral7B
Number Token $p$-value Comment
1 CV 0.000000241 Must start a word
2 ember 0.00000199
3 ¡Korean character¿ 0.00000224
4 ĭ 0.00000770
5 rise 0.0000283 Must start a word
6 poured 0.0000305 Must start a word
7 aimed 0.0000364 Must start a word
8 allocate 0.0000553 Must start a word
9 UINT 0.0000684
10 monitor 0.0000767 Must start a word
11 ¡Cyrillic¿ 0.000105
[…]
13 ¡Chinese character¿ 0.000133
14 Lat 0.000139 Must start a word
15 Vel 0.000148
16 challenges 0.000178 Must start a word
17 ¡Chinese character¿ 0.000200
18 Od 0.000222 Must start a word
19 neys 0.000244
20 selon 0.000244 Must start a word
21 measures 0.000268 Must start a word
22 assumes 0.000321 Must start a word
23 Et 0.000338 Must start a word
24 Rot 0.000438
25 ¡Chinese character¿ 0.000538
26 olutely 0.000540
27 HO 0.000567
28 Solutions 0.000598 Must start a word
29 * 0.000658
30 Condition 0.000665
31 zt 0.000677
32 iced 0.000696
33 URL 0.000722
34 SEO 0.000728 Must start a word
35 weigh 0.000728 Must start a word
36 änge 0.000765
37 Vic 0.000777 Must start a word
38 apparent 0.000780 Must start a word
39 ¡Chinese character¿ 0.000896
40 ¡0xE3¿ 0.000921
Table 8. Rejections of the manifold hypothesis for Pythia6.9B (1 of 2)
Number Token $p$-value Comment
1 psychotic 0.000000161 Must start a word
2 \\]( 0.00000147
3 1990 0.00000231
4 tiene 0.00000282 Must start a word
5 ucky 0.0000134
6 Hebrew 0.0000141 Must start a word
7 mess 0.0000184
8 reasonable 0.0000300 Must start a word
9 collector 0.0000311 Must start a word
10 embodiments 0.0000428 Must start a word
11 Inhibition 0.0000593 Must start a word
12 Hers 0.0000691 Must start a word
13 odge 0.000120
14 áĢIJ 0.000144
15 ÏĢο 0.000165
16 198 0.000170 Must start a word
17 scare 0.000192 Must start a word
18 ]{}\\_[ 0.000202
19 LIMITED 0.000212 Must start a word
20 hop 0.000232 Must start a word
21 BIO 0.000246
22 raison 0.000252 Must start a word
23 inside 0.000254
24 st 0.000261
25 waving 0.000286 Must start a word
26 Austria 0.000287 Must start a word
27 ny 0.000313
Table 9. Rejections of the manifold hypothesis for Pythia6.9B (2 of 2)
Number Token $p$-value Comment
28 bags 0.000326
29 medium 0.000372
30 incent 0.000378 Must start a word
31 carboh 0.000407 Must start a word
32 secret 0.000451 Must start a word
33 ¡Unicode characters¿ 0.000468
34 ÏĢε 0.000519 Must start a word
35 120 0.000536
36 wastes 0.000565 Must start a word
37 ressor 0.000570
38 Lip 0.000597 Must start a word
39 preview 0.000650
40 redund 0.000653 Must start a word
41 ridic 0.000673 Must start a word
42 Ramirez 0.000698 Must start a word
43 omatic 0.000707
44 Gloria 0.000733 Must start a word
45 eton 0.000780
46 chairman 0.000789 Must start a word
47 Supreme 0.000859 Must start a word
48 grazing 0.000890 Must start a word
49 activation 0.000892 Must start a word
50 imming 0.000914
51 radiological 0.000931 Must start a word
52 branc 0.000941 Must start a word
53 \$\\—\\ 0.000942 Must start a word
54 DEFAULT 0.000992
See Tables 49.